OpenStudy (anonymous):

A circle has the equation x^2+y^2-3x-4y+4=0 Find the center in terms of (h,k) and any x or y intercepts.

OpenStudy (anonymous):

Hint: Complete the square on each of the variables \(x, y\)

OpenStudy (anonymous):

Yeah I can get that done, can't seem to find the correct intercepts for some reason.

OpenStudy (anonymous):

For the \(x\)-intercepts, we simply plug in \(y=0\), so: \[ x^2+y^2-3x-4y+4=0\implies\\ x^2-3x+4=0\implies x\in\mathbb{C}\setminus\mathbb{R} \]So, there are no \(x\)-intercepts. Try doing the process above, for \(y\).

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

Sure thing.